OpenStudy (anonymous):

I need help on the first one, and I need someone to check the two I did on my own: 1. Given the equation -4squrt(x-3)=12, solve for x and identify if it is an extraneous solution. x = 0, solution is not extraneous x = 0, solution is extraneous x = 12, solution is not extraneous x = 12, solution is extraneous

OpenStudy (anonymous):

don't even start

OpenStudy (anonymous):

2. Given the equation \[\sqrt{8X+1}=5\] x = One fourth, solution is not extraneous x = One fourth, solution is extraneous x = 3, solution is not extraneous x = 3, solution is extraneous (my ans x = One fourth, solution is not extraneous x = One fourth, solution is extraneous x = 3, solution is not extraneous x = 3, solution is extraneous were)

OpenStudy (anonymous):

\[-4\sqrt{x-3}=12\] has no solution because the left hand side is negative and the right hand side is positive

OpenStudy (anonymous):

Oh so it would be x = 0, solution is not extraneous

OpenStudy (anonymous):

no, it would be no solution because the left hand side is negative and the right hand side is positive

OpenStudy (anonymous):

The second on I did, I got x = 3, solution is extraneous I just want to know if I'm right

OpenStudy (anonymous):

Oh...no solution isn't a option

OpenStudy (anonymous):

and I need to see if im right about the one above and this one: \[\sqrt{x+9}-4=1\] x = 16, solution is not extraneous x = 16, solution is extraneous (MY ANSWERE) x = 34, solution is not extraneous x = 34, solution is extraneous

OpenStudy (anonymous):

2. Given the equation \[\sqrt{8X+1}=5\] x = One fourth, solution is not extraneous x = One fourth, solution is extraneous x = 3, solution is not extraneous x = 3, solution is extraneous (my answer)